### v2

# Mixture-of-RAG: Integrating Text and Tables with Large Language Models

Abstract
Large language models (LLMs) achieve optimal utility when their responses are grounded in external knowledge sources. However, real-world documents, such as annual reports, scientific papers, and clinical guidelines, frequently combine extensive narrative content with complex, hierarchically structured tables. While existing retrieval-augmented generation (RAG) systems effectively integrate LLMs’ generative capabilities with external retrieval-based information, their performance significantly deteriorates especially processing such heterogeneous text-table hierarchies. To address this limitation, we formalize the task of Heterogeneous Document RAG, which requires joint retrieval and reasoning across textual and hierarchical tabular data. We propose MixRAG, a novel three-stage framework: (i) hierarchy row-and-column-level (H-RCL) representation that preserves hierarchical structure and heterogeneous relationship, (ii) an ensemble retriever with LLM-based reranking for evidence alignment, and (iii) multi-step reasoning decomposition via a RECAP prompt strategy. To bridge the gap in available data for this domain, we release a large-scale dataset DocRAGLib, a 2k-document corpus paired with automatically aligned text-table summaries and gold document annotations. The comprehensive experiment results demonstrate that MixRAG boosts top-1 retrieval by 46% over strong text-only, table-only, and naive-mixture baselines, establishing new state-of-the-art performance for mixed-modality document grounding.
1 Introduction
With the rapid development of large language models (LLMs) [20; 14], the reliability of their response substantially depends on the extent to which they can be grounded in verifiable evidence. Retrieval-Augmented Generation (RAG) has emerged as an effective strategy that combines the generative power of LLMs with external retrieval-based information. Recent research has extended the application of RAG to various scenarios, including knowledge graph [24; 17], database [2], and multi-modal data [15; 32]. When the external source is plain text or simple tables, RAG systems retain precision and fluency. However, real-world data, including financial statements, clinical guidelines, research articles, and governmental reports, all interleave complex hierarchical tables whose rows and columns nest multiple semantic layers. The crucial metrics numbers, such as earnings per share, dosage limits, and experimental results, are embedded within these tables, and the relationships among those numbers are defined by the headers that span multiple rows or columns. Therefore, the research on the Heterogeneous Document RAG remains nascent. The Heterogeneous Document RAG aims to efficiently retrieve and synthesize relevant information from large, complex documents by integrating textual and hierarchical tabular data with the power of LLMs for more comprehensive generation.
Figure 1: An Example of Heterogeneous Document RAG Task
The Heterogeneous Document RAG task involves retrieving the most relevant documents from extensive document repositories based on a specific question and generating the corresponding answer. However, it is challenging due to the complexity and scale of real-world repositories, which often contain numerous indistinguishable documents with intricate structures. For example, Statistics Canada offers thousands of statistical reports across subjects, including 1,761 on "Labour" and 1,132 on "Population and demography", with semantically similar content often combining textual descriptions and hierarchical tables.
Some novel RAG techniques  [13; 3; 21; 12] have been explored to improve retrieval accuracy by enhancing RAG frameworks. While effective in text-based tasks, these methods often fail to capture the critical table information within retrieved documents. Other table-focused RAG techniques [2; 18; 10; 17] lack seamless integration of table data with text, resulting in unaligned representations and difficulty handling complex hierarchical structures. As shown in Figure 1, when retrieval relies on the hierarchical structure of a table, such methods often break down the complex structure, leading to information loss, i.e., disruption of critical dependencies (e.g., 2012-2014 trends), and a reduced ability to capture critical relationships within the data. Additionally, while well-established document-based QA methods [27; 19; 26] have demonstrated relatively high accuracy, they still struggle with complex tasks involving documents with both long text and hierarchical tables [31]. (More related work attached in Appendix I.) Overall, although existing methods have provided preliminary solutions for Heterogeneous Document RAG, achieving more precise information retrieval and integration still faces numerous challenges: (i) Data Scarcity. The publicly available benchmarks lack a cleaned, integrated corpus pairing narrative text with hierarchical tables. (ii) Structured Representation. The unified representation must preserve row-column hierarchies while remaining consistent with text. As shown in Figure 1, the key terms "Total Buy" and "Revenues of Electric Sales" are strongly linked to the hierarchical structure within the heterogeneous document, and the information contained within tables is dense, highlighting the challenge of representing structured data in such documents. (iii) Retrieval Precision. The traditional semantic similarity-based retrieval methods often return text chunks that, while semantically related to the question, may not provide valid evidence for reasoning. For example, as illustrated in Figure 1, existing methods may retrieve documents containing the exact year ’2014’ from the query, while missing the relevant historical data from 2012 and 2013 that the query actually seeks. (iv) Reasoning Accuracy. Effective retrieval requires hierarchical alignment between unstructured text and structured tabular data, necessitating the ability to process and reason across heterogeneous documents accurately (e.g., deriving 2012-2013 context from a 2014 query and multi-step computation).
To address the above challenges in Heterogeneous Document RAG, we propose the DocRAGLib dataset and MixRAG framework. Specifically, DocRAGLib consists of a document repository with 2,178 documents, each of which combines text and hierarchical tables, as well as 4,468 QA pairs. DocRAGLib is collected from reliable sources and cleaned to ensure consistency across document types and formats. To the best of our knowledge, DocRAGLib is the first dataset that bridges the gap of heterogeneous documents. MixRAG framework consists of three modules: Heterogeneous Document Representation Module, Cross-Modal Retrieval Module, and Multi-Step Reasoning Module. Specifically, the Heterogeneous Document Representation Module employs a hierarchical row-and-column-level (H-RCL) table summarization method to capture the structure and content of complex tables. It generates representations that preserve table structures and are optimized for retrieval, addressing the challenge of effectively representing table information in heterogeneous documents. The Cross-Modal Retrieval Module overcomes the limitations of relying solely on semantic similarity retrieval by adopting a two-stage approach. First, the ensemble retrieval stage combines BM25 with the semantic understanding of embedding retrieval to filter candidate documents. Second, the LLM-based reranking stage utilizes the contextual reasoning ability of LLMs to identify the most relevant document, improving the accuracy and comprehensiveness of retrieval. The Multi-Step Reasoning Module introduces the RECAP prompt strategy, which decomposes complex reasoning into sub-tasks and leverages external calculators to manage mathematical operations during the reasoning process. MixRAG significantly improves answer accuracy and effectively addresses the challenges of multi-step reasoning and complex calculations in heterogeneous documents.
In summary, our main contributions are as follows:
  • We propose a new large-scale dataset, DocRAGLib, which provides a high-quality benchmark for heterogeneous information retrieval tasks.
  • We design the MixRAG framework, which consists of three components: the Heterogeneous Document Representation Module, the Cross-Modal Retrieval Module, and the Multi-Step Reasoning Module.
  • We conduct extensive experiments to validate the superior performance of the MixRAG framework over baseline methods, with ablations verifying the effectiveness of each module and scalability analysis showing the framework’s robustness across different corpus sizes.
2 Preliminary
In this section, we formally define the Heterogeneous Document RAG task and detail the collection, processing, and cleaning procedures for our dataset, DocRAGLib.
2.1 Task Definition
In this section, we formally define key concepts and our Heterogeneous Document RAG task.
Document Corpus. A comprehensive collection of documents $\mathbf{C}=\{D_{1},D_{2},\dots,D_{|\mathbf{C}|}\}$ where each document $D_{i}\in\mathbf{C}$ consists of multiple text segments $\mathbf{P}=\{P_{1},P_{2},\dots,P_{|\mathbf{P}|}\}$ and several complex hierarchical tables $\mathbf{T}=\{T_{1},T_{2},\dots,T_{|\mathbf{T}|}\}$, referred to as heterogeneous document. Each hierarchical table is defined as a tuple $T=\{{H}_{t},{H}_{l},d\}$, where ${H}_{t}$ denotes the top headers in hierarchical structure of table $T$, ${H}_{l}$ denotes the left headers in hierarchical structure of table $T$ (example of hierarchical headers shown in Figure 1), and $d$ denotes the data cells of table $T$.
Question. A question $Q$ posed by the user, for which there exists a unique document of highest relevance $D^{*}\in\mathbf{C}$ in the corpus.
Given a question $Q$ and a document corpus $\mathbf{C}$, our task is to retrieve the most relevant document $D^{*}$ and generate a comprehensive answer $A$ based on the information contained within $D^{*}$. Formally, the Heterogeneous Document RAG task can be described by the following function:
$A=\mathcal{F}(Q,\mathbf{C})=\text{Inference}\left(Q,D^{*}\right),\text{where }D^{*}=\text{Retrieve}(Q,\mathbf{C})$ (1)
where $\text{Retrieve}(Q,\mathbf{C})$ is the function that retrieves the most relevant document $D^{*}\in\mathbf{C}$ based on the question $Q$, and $\text{Inference}(Q,D^{*})$ is the generation function that returns an accurate answer $A$ by leveraging the information contained within $D^{*}$.
2.2 Dataset
To bridge the gap of heterogeneous data availability, we construct the DocRAGLib dataset based on existing relevant datasets  [30; 9] and data crawled from the web. DocRAGLib contains 2,178 documents, combining text and table content, and includes 4,468 QA pairs designed based on the content of these documents. Existing datasets are predominantly tailored for single-document RAG tasks, limiting their applicability to scenarios requiring the processing of large-scale and complex document corpora. DocRAGLib aims to address this limitation by providing a comprehensive dataset specifically designed for multi-document RAG tasks. The details of Data Collection, Data Cleaning, and Data Statistics are elaborated in Appendix A.
3 Methodology
To address this problem, we introduce MixRAG, a framework designed for the Heterogeneous Document RAG Task.
Figure 2: The Overview of our MixRAG Framework
As depicted in Figure 2, MixRAG comprises three modules: Heterogeneous Document Representation Module, Cross-Modal Retrieval Module, and Multi-Step Reasoning Module. The following sections show a detailed introduction to the three modules.
3.1 Heterogeneous Document Representation Module
The Heterogeneous Document Representation Module focuses on generating effective representations of heterogeneous documents, capturing both lengthy text and complex hierarchical tables to optimize retrieval performance. A traditional approach, the Table Level Summary, involves providing the table’s schema to the LLM and summarizing it. However, it oversimplifies the table, leading to the loss of key information and ignoring important structural details. For effective representation, the Heterogeneous Document Representation Module first introduces a row-and-column-level(RCL) table summary to represent tables. Then, leveraging the hierarchical structure of the tables, we extend this RCL table summary to capture the nested relationships and dependencies within the table data, ensuring a more comprehensive representation (see Appendix B for examples and comparison of different representation strategies).
To better understand the proposed RCL table summary, we introduce some basic concepts related to table structure and paths.
For a hierarchical table, as we defined before, table $T=\{H_{t},H_{l},d\}$, where $H_{t}$ represents the top headers, $H_{l}$ represents the left headers, and $d$ represents the data cells. Specifically, we can define:
$\left\{\begin{aligned} H_{t}&=\{H_{t}^{r}\mid r=1,\dots,R_{t}\}\\ H_{l}&=\{H_{l}^{r}\mid r=1,\dots,R_{l}\}\end{aligned}\right$. (2)
, where $R_{t}$ denotes the number of hierarchical levels in the top headers, and $R_{l}$ denotes the number of hierarchical levels in the left headers.
For each hierarchical level in the top and left headers, we define:
[…]
, where $h_{t}^{r}(j)$ represents the header at level $r$ in the top header hierarchy, while $h_{l}^{r}(i)$ represents the header at level $r$ in the left header hierarchy. As shown in Figure 3, these headers form multi-level paths used to locate specific data cells within the table. To pinpoint a particular data cell $d_{ij}$, we define the paths from both the left and top headers, $P_{l}(i)$ and $P_{t}(j)$, respectively, which together uniquely identify the location of the data cell:
$\left\{\begin{aligned} P_{l}(i)&=h_{l}^{1}(i_{1})\to h_{l}^{2}(i_{2})\to\dots\to h_{l}^{R_{l}}(i_{R_{l}})\\ P_{t}(j)&=h_{t}^{1}(j_{1})\to h_{t}^{2}(j_{2})\to\dots\to h_{t}^{R_{t}}(j_{R_{t}})\end{aligned}\right$. (4)
These paths capture the hierarchical relationships between the cells in the table, allowing us to pinpoint any data cell $d_{ij}$ by referencing the corresponding paths in the left and top headers.
Below, we provide a detailed formal definition of the proposed RCL table summary, outlining its structure and how it incorporates hierarchical relationships for enhanced table representation.
3.1.1 General RCL Table Summary
In the case of a general table, where multi-level hierarchies and complex relationships within the table are not considered, we summarize each row and column independently. By disregarding the hierarchical structure, the table is reduced to a flat, 1-level representation, where both the top and left headers are flattened.
Formally, when the hierarchical structure is ignored, $H_{t}$ and $H_{l}$ are reduced to:
$\left\{\begin{aligned} H_{t}&=\{h_{t}(1),h_{t}(2),\dots,h_{t}(n)\mid n\text{ is \# columns}\}\\ H_{l}&=\{h_{l}(1),h_{l}(2),\dots,h_{l}(m)\mid m\text{ is \# rows}\}\end{aligned}\right$. (5)
Similarly, the paths simplify to $P_{l}(i)=h_{l}(i)$ and $P_{t}(j)=h_{t}(j)$. In this generalized case, each row and column is represented by a left and top header, respectively. To summarize the table at the general row and column level, we flatten the table’s hierarchical structure and treat it as a single level. Specifically, for each row and column, the general RCL table summary can be formulated as:
$\left\{\begin{aligned} S_{\text{row}}\{h_{l}(i)\}&=\mathcal{G}(h_{l}(i),\bigcup_{j=1}^{n}h_{t}(j))\\ S_{\text{col}}\{h_{t}(j)\}&=\mathcal{G}(h_{t}(j),\bigcup_{i=1}^{m}h_{l}(i))\end{aligned}\right$. (6)
, where $\mathcal{G}$ denotes the operation that generates the general RCL summary for both rows and columns. This operation condenses the information in each row and column, simplifying the table’s hierarchical structure for retrieval.
3.1.2 H-RCL Table Summary
Hierarchical tables contain multiple levels of headers that form a complex hierarchy, enriching the data structure. In this context, we extend the General RCL table summary to handle multi-level structures, which capture the intricate dependencies between the table’s headers and data more effectively.
We propose the H-RCL table summary that preserves the nested relationships between rows and columns by leveraging the hierarchical paths of both top and left headers, as shown in Eq. 3 and Eq. 4. In the following, we provide a formal definition of the hierarchical row and column summaries:
$\left\{\begin{aligned} S_{\text{row}}(h_{l}^{r}(i))&=\mathcal{G}(P_{l}^{r}(i),\bigcup_{j=1}^{n}P_{t}^{R_{t}}(j))\\ S_{\text{col}}(h_{t}^{r}(j))&=\mathcal{G}(P_{t}^{r}(j),\bigcup_{i=1}^{m}P_{l}^{R_{l}}(i))\end{aligned}\right$. (7)
The row-level summaries $S_{\text{row}}$ capture the dependencies within the left headers, while the column-level summaries $S_{\text{col}}$ reflect the relationships within the top headers. As shown in Figure 3, the $S_{\text{row}}$ aggregates information from both the left and top headers at their respective hierarchical levels. After aggregation, the information is provided to the LLM for refinement to generate the final summary while preserving the complex relationships within the table’s structure. These summaries jointly represent the hierarchical structure of the table. The overall table summary is defined as:
$S_{\text{table}}=\bigcup_{i,r}S_{\text{row}}(h_{l}^{r}(i))\cup\bigcup_{j,r}S_{\text{col}}(h_{t}^{r}(j))$ (8)
By combining both row and column summaries, the table summary $S_{\text{table}}$ offers a comprehensive representation of the hierarchical table, effectively maintaining the nested structure of the data.
3.1.3 Passage Processing
In addition to processing tables, we also address the handling of text passages within the heterogeneous document. In this module, we focus on processing the passage at the sentence level, dividing the passage into individual sentences, and constructing the corpus. We segment the passage at the sentence level to enable more effective handling of textual data, ensuring that relevant information is preserved and structured for retrieval and reasoning.
Figure 3: The Path and Hierarchical Levels in the Table. The H-RCL Table Summary Relies on the Structure and Paths of Hierarchical Table
3.2 Cross-Modal Retrieval Module
The Cross-Modal Retrieval Module is designed to identify the most relevant document $D^{*}$ from the document corpus for a given question. It consists of two key stages: Ensemble retrieval and LLM-based retrieval. The ensemble retrieval stage combines BM25 and embedding-based retrieval to balance keyword matching with semantic understanding. BM25 captures exact term relevance, while embedding-based retrieval ensures deeper semantic alignment, enabling the retrieval of a comprehensive set of top-$k$ candidate documents. Specifically, we utilize two different types of corpora for the ensemble retrieval stage, tailored to align with the underlying mechanisms of embedding-based and BM25-based approaches. The ensemble retrieval stage conducts a coarse-grained filtering of documents across the entire document corpus, ensuring that the candidate documents provided to the LLM remain within its maximum input length constraints. The LLM-based retrieval performs fine-grained logical reasoning tailored to the specific question and selected candidate documents, leveraging the contextual reasoning capabilities of the LLM to prioritize documents most relevant to the question. This approach ensures high precision in the final output.
The Ensemble Retrieval Stage. The ensemble retrieval stage is designed to generate a robust set of top-$k$ candidate documents by combining the complementary strengths of BM25 and embedding-based retrieval methods. Below, we detail the three key components of this stage: BM25 retrieval, embedding-based retrieval, and candidate merging. (1) BM25 Retrieval. We use BM25 to rank chunks from the BM25 corpus based on their relevance to a given question $Q$. Each chunk corresponds to a specific section of its document. BM25 emphasizes exact keyword matches and considers term frequency within a chunk while normalizing for document length to ensure fair comparisons. By prioritizing terms that are frequent in the chunk but rare across the corpus, BM25 effectively retrieves the top-$n$ candidate documents that are most aligned with the question. As the RCL method may over-aggregate keywords and BM25’s common-term bias, we use the corpus constructed with the table summary method for retrieval at this stage. (2) Embedding-based Retrieval. Simultaneously, embedding-based retrieval identifies the top-$m$ candidate documents from the embedding corpus by computing semantic similarity between the question $Q$ and each chunk. Both the question and chunks are represented as dense vectors generated by the embedding model. The embedding-based retrieval method captures semantic nuances and retrieves chunks that may not share explicit terms with the question but are contextually relevant. In this retrieval stage, we utilize the corpus constructed using the H-RCL method, ensuring that the embedding model can more precisely capture information semantically similar to the question within hierarchical tables. (3) Candidate Merging. After retrieving the top-$n$ candidates from BM25 and top-$m$ from embedding-based retrieval, the two sets are merged into a final top-$k$ set with duplicates removed. This approach combines the strengths of both methods, ensuring a diverse and semantically rich candidate pool.
The ensemble approach combines the strengths of BM25 and embedding-based retrieval to create a diverse candidate document set. BM25 ensures accurate keyword-based matches, while embedding-based retrieval adds semantic depth and flexibility. By merging, the ensemble retrieval stage provides high recall and relevance, laying a strong foundation for the subsequent LLM-based retrieval stage.
The LLM-based Retrieval Stage. After the preceding Ensemble Retrieval Stage, the Cross-Modal Retrieval Module extracts the top-$k$ candidate documents from the original document corpus that are most similar to the question. To further filter out the target document that is most relevant to the question, we introduce LLM as an expert in similarity analysis. Specifically, each question alongside its corresponding candidate documents is presented to the LLM, which is tasked with determining the single document most relevant to the question. However, since each document is typically lengthy, simultaneously inputting the top-$k$ documents into the LLM often exceeds the model’s maximum input token limit. Moreover, the target document may contain a large amount of irrelevant information, which can reduce the LLM’s accuracy when processing the entire content. To mitigate these challenges, we design the input space by setting a variable filtering threshold, $\Theta$, and selecting the top-$\Theta$ information chunks most similar to the question from all documents. In this process, an external function ensures that $\Theta$ is set to its maximum value, provided that the token length of these chunks does not exceed the model’s input limit. These selected chunks are then consolidated into a single collection:
$S=\text{SelectTop}(Q,D,\Theta)$ (9)
where $Q$ represents the original question. $D$ denotes the set of all documents. $\Theta$ is the filtering threshold. $S$ represents the set of selected information chunks.
Subsequently, for each document $D_{i}$ in the top-$k$ documents, only the information belonging to this collection is retained, and the filtered information is recombined into a "new document" $D^{\prime}_{i}$ to replace the original document:
$D^{\prime}_{i}=\text{Recombine}(S,D_{i})$ (10)
Finally, we input the question content $Q$ and its corresponding top-$k$ new documents $\{D^{\prime}_{1},D^{\prime}_{2},\ldots,D^{\prime}_{k}\}$ into the LLM for similarity analysis, thereby selecting the most relevant document $D^{*}$ among the top-$k$ documents.
3.3 Multi-Step Reasoning Module
We execute QA tasks as an instance based on the most relevant document $D^{*}$ provided by the Cross-Modal Retrieval Module. Although LLMs demonstrate strong performance in natural language understanding, performing QA on heterogeneous documents remains a challenging task. To more accurately extract relevant information from heterogeneous documents and leverage the information to answer the questions, we introduce a novel prompting strategy for QA tasks with MixRAG called RECAP (Restate, Extract, Compute, Answer, Present), inspired by prompt strategies such as CoT [25] and EEDP [19] that involve step-by-step analytical reasoning, as well as PoT [6] which generates executable programs to solve complex mathematical operations.
Formally, as defined in Section 2.1, the Multi-Step Reasoning Module is represented as $A=\text{Inference}(Q,D^{*})$, where $A$ is the final answer produced by the RECAP strategy.
The specific workflow of RECAP is as follows: (1) Restate the Question: Clearly restate the requirements of the question and analyze the question type; (2) Extract Relevant Data: Extract and list all the information from the heterogeneous document that is directly relevant to the question; (3) Perform Logical Reasoning: The LLM is required to perform step-by-step reasoning to obtain the final answer, and the LLM should also provide the complete calculation formula used to derive the final result (if the question is not calculation-based, the result is provided directly); (4) Provide Final Answer: For the answer derived by the LLM through reasoning and the other answer obtained by the external calculator based on the provided formula, we employ a rule-based approach to select the more appropriate answer from the two, depending on the question type, as the final answer. Unlike methods like Self-Consistency [23], which improve accuracy by generating multiple reasoning paths, RECAP adopts a single-turn dialogue strategy, with its logical reasoning and formula provision originating from the same reasoning path, offering advantages in token efficiency for the LLM. The comparison between RECAP and various QA inference baselines is given in Appendix  G.
4 Experiment
4.1 Experimental Setup
To provide a comprehensive evaluation of MixRAG, we conduct experiments comparing its performance with relevant baseline methods. Detailed descriptions of these baselines are provided in Appendix C. We adopt HiT@K and Exact Match (EM) as evaluation metrics for retrieval and QA, respectively. Further details are available in Appendix D. We employ GPT-4o throughout our pipeline, including heterogeneous document representation, LLM-based retrieval, and RECAP execution. For embedding-based retrieval, we use OpenAI’s text-embedding-3-large model as the encoder. According to Appendix F, we set the parameters of MixRAG to $n=40$ and $m=60$ for ensemble retrieval. For the QA instance RECAP, the number of examples for all prompt strategies is uniformly set to 1. We evaluate the RECAP instance using GPT-4o, GPT-4o mini, and three open-source large language models: Qwen2.5-32B-Instruct, Mistral-Nemo-Instruct-2407, and Llama-3.1-8B-Instruct. Detailed evaluation results are presented in Appendix G.
4.2 Comparison with Baselines
In this section, we present a comprehensive analysis of MixRAG compared with several strong retrieval-augmented baselines, across the DocRAGLib dataset to evaluate both retrieval performance (HiT@K) and downstream QA accuracy (EM).
Table 1: Comparison of MixRAG Performance with Various Baselines
HiT@K EM*
Method K=1 K=3 K=5 K=10
Standard RAG 0.0159 0.0339 0.0538 0.1255 0.0092
LangChain 0.2390 0.4104 0.4821 0.5219 0.1301
Self-RAG 0.2829 0.4502 0.5000 0.5598 0.2141
Table Retrieval 0.3705 0.5299 0.5996 0.6494 0.2541
MixRAG 0.5410 0.7244 0.7603 0.8689 0.3228
*Note: EM scores are based on HiT@1 retrieval using RECAP.
Retrieval Performance Comparison with Baselines. Table 1 presents the retrieval performance of MixRAG compared to several baselines. Across all HiT@K metrics (K=1, 3, 5, 10), MixRAG consistently achieves the best results, reaching a HiT@1 score of 0.541, significantly outperforming the strongest baseline by 46%. The fine-grained row-and-column-level summarization, combined with the two-stage retrieval strategy, achieves substantial improvements in retrieval accuracy for complex heterogeneous documents.
Document-QA Accuracy Comparison with Baselines. Using RECAP as the QA instance module for each retrieval method, we evaluate document-level QA accuracy based on EM scores. According to the EM results in Table 1, MixRAG also achieves the highest QA accuracy, with an EM score of 0.3228, which corresponds to a 27.1% improvement over the second-best baseline, Table Retrieval. These results highlight the effectiveness of our retrieval and reasoning framework in improving answer accuracy for heterogeneous documents retrieval in downstream QA tasks.
4.3 Ablation Study
To systematically investigate the effectiveness of each key component in MixRAG, we conduct comprehensive ablation studies based on the DocRAGLib dataset. The studies include variations in the heterogeneous document representation strategies and MixRAG modules.
Impact of Heterogeneous Document Representation Strategy in MixRAG. To evaluate the impact of different table summarization methods in heterogeneous document representation module, we compare three variants: Table Level Summary, General RCL Summary, and H-RCL Summary. As shown in Table 2, the H-RCL Table Summary strategy achieves the best performance among all heterogeneous document representation strategies. The General RCL Summary ranks second, showing clear gains over the traditional Table Level Summary by providing more explicit representations of rows and columns. H-RCL improves HiT@1 by 47% and EM by 28% compared to the Table Level Summary, demonstrating the importance of capturing multi-level table structures. The General RCL strategy provides a clearer representation of rows and columns, reducing ambiguity and improving retrieval accuracy, which in turn enhances QA performance by enabling more precise document matching. Building on this, the H-RCL strategy further captures multi-level table structures and incorporates hierarchical path information, significantly boosting both retrieval accuracy and the model’s ability to perform complex reasoning over structured data. Overall, these results demonstrate the effectiveness of the H-RCL design in improving both retrieval and reasoning capabilities for heterogeneous documents.
Table 2: Performance Evaluation of MixRAG with Different Table Summary Level
heterogeneous document representation HiT@K EM*
Table Summary Level K=1 K=3 K=5 K=10
Table Level Summary 0.3627 0.4805 0.5963 0.7428 0.2383
General RCL Summary 0.3984 0.5502 0.6803 0.7981 0.2625
H-RCL Summary 0.5410 0.7244 0.7603 0.8689 0.3228
*Note: EM scores are based on HiT@1 retrieval using RECAP.
Impact of Retrieval and QA Module. To investigate the effectiveness of each key component in MixRAG, we conduct the ablation study using DocRAGLib dataset. Table 4.3 shows the results of the Retrieval and QA module ablation study. Replacing Ensemble Retrieval with either Embedding or BM25 significantly reduces HiT@1, with Embedding achieving better results due to its stronger semantic alignment capabilities. Further removing LLM-based retrieval leads to an additional decline of 0.19 and 0.20 in HiT@1 for Embedding and BM25, respectively, confirming the advantage of the two-stage retrieval framework over single-stage methods. Moreover, disabling external calculation in RECAP results in a noticeable drop in EM, underscoring the importance of precise formula-based reasoning for improving answer accuracy. Overall, these results validate the effectiveness of each component within MixRAG in enhancing both retrieval precision and answer accuracy.
Table 3: Ablation Study Result of MixRAG
HiT@1 EM
Ours (MixRAG) 0.5410 0.3238
- w/o Ensemble Retrieval
only Embedding 0.4734 0.2777
only BM25 0.3545 0.2100
- w/o Two-stage Retrieval*
only Embedding 0.2787 0.2131
only BM25 0.1557 0.1311
- w/o Calc of RECAP 0.5410 0.2838
*Note: w/o both Ensemble Retrieval and LLM-based Retrieval.
Figure 4: Scalability Analysis of Retrieval Performance across Different Corpus Sizes.
We evaluate the scalability of MixRAG by comparing its retrieval performance across different document corpus sizes. As shown in Figure 4, MixRAG consistently maintains the highest HiT@K scores (K=1, 3, 5, 10) across all corpus scales, demonstrating its ability to preserve retrieval precision even as the corpus size increases. While baseline methods experience significant drops in performance as the document corpus increases, MixRAG exhibits the slowest rate of degradation, particularly in HiT@10, where its performance remains notably stable even with a larger document corpus. The retrieval stability reflects the scalability of MixRAG in handling growing document complexity. These results confirm that MixRAG effectively maintains retrieval precision even with larger and more complex document sets, showcasing its superior scalability compared to existing methods.
5 Conclusion and Limitations
Conclusion. In this paper, we propose MixRAG, a framework designed for Heterogeneous Document RAG. To support this task, we introduce DocRAGLib, a benchmark dataset containing over 2000 heterogeneous documents and over 4000 QA pairs. MixRAG integrates hierarchical table summarization, cross-modal retrieval with LLM-based reranking, and RECAP-guided multi-step reasoning to support RAG over heterogeneous documents. Extensive experiments show that MixRAG outperforms existing baselines in retrieval precision and QA accuracy, establishing new state-of-the-art results for mixed-modality document grounding.
Limitations. Our Heterogeneous Document RAG task setup selects the top-1 retrieved document, which may limit performance on tasks requiring evidence aggregation from multiple document sources. In the future, extending the framework to incorporate multi-document retrieval and integration could better capture distributed evidence and support complex scenarios. In addition, MixRAG specializes in heterogeneous documents integration. It does not handle visual modalities (images, figures, or schematics) commonly found in real-world documents. Incorporating visual modalities into the RAG framework presents a promising direction for future research.
Appendix A Dataset
A.1 Data Collection
The documents of DocRAGLib are derived from two sources: the existing public single-document QA dataset and web-scraped text-table heterogeneous data. First, we collect information from the publicly available single-document QA dataset MultiHiertt [30], which is designed for single-document QA tasks. Each document in MultiHiertt comprises textual content and hierarchical tables, with an average of 3.89 tables per document. In addition, we scrape contextual information related to the tables by utilizing the webpage links provided in another available single-table QA dataset HiTab [9], and then reconstruct heterogeneous documents containing both text and tables.
A.2 Data Cleaning
The raw data consists of heterogeneous documents and QA pairs. This section outlines the data cleaning process, focusing on the strategies to refine both documents and QA pairs.
For document refinement, we reduce noise in table contexts extracted from web pages by addressing issues such as irregular delimiters and unrelated sentences. Noise removal is performed through manual cleaning, which involves (1) reformatting inconsistent delimiters and (2) removing irrelevant text segments.
Figure 5: Pipeline for ambiguity resolution during QA pair construction
To ensure clarity in the QA pairs, we further resolve ambiguities arising when multiple documents could potentially answer the same question. The process begins with rule-based filtering to identify candidate documents. Each candidate is then evaluated for relevance by pairing it with the question and inputting the pair into GPT-3.5 Turbo and GPT-4o. Relevance judgments are tracked iteratively, and if the number of remaining documents falls below a predefined threshold, the question is considered unambiguous and the corresponding QA pair is retained. As shown in Figure 5, this process systematically eliminates ambiguous questions and ensures the overall quality of the dataset.
A.3 Data Statistics
After the data collection and cleaning process, 2,178 heterogeneous documents and 4,468 QA pairs are obtained. To enable further exploration of pre-trained model-based methods for the Heterogeneous Document RAG task, the 4,468 QA pairs are split into three subsets: train (2,990 QA pairs), dev (502 QA pairs), and test (976 QA pairs). To ensure consistency, the 2,178 heterogeneous documents are merged into a single collection, without further division.
As shown in Table 4, statistical analysis of the DocRAGLib dataset reveals that: (1) The average textual length of each document is 1,453.05 words, and the average number of tables is 3.85, with 88.71% of documents containing at least two tables. (2) The QA pairs are categorized by evidence source: 9.02% of questions rely solely on text, 50.78% on tables, and 40.20% on both. Additionally, we analyze the numerical scale and precision of answers in the DocRAGLib dataset. Specifically, 19.36% of answers contain values exceeding 10,000, 4.05% exceed 100,000, and 34.29% are decimals with more than three significant digits. These statistics highlight the complexity and diversity of the dataset, offering a challenging benchmark for future research.
Table 4: Statistical Analysis of the DocRAGLib Dataset
Indicator Value
Total Document Number 2,178
    Avg. Word Number per Document 1,453.05
    Avg. Table Number per Document 3.85
    Documents with >2 Tables 1,932 (88.71%)
Total QA Pairs Number 4,468
    Train/Dev/Test 2,990/502/976
Classification by Evidence Source
    Text Only 403 (9.02%)
    Table Only 2,269 (50.78%)
    Both Text and Table 1,796 (40.20%)
Answer Characteristics
    Decimal Places $\geq$3 1,532 (34.29%)
    Numerical Values >10,000 865 (19.36%)
    Numerical Values >100,000 181 (4.05%)
Appendix B Different Table Representation Strategies
To understand how summarization granularity affects information quality and retrieval, we compare three table summarization strategies: Table-Level, Row Column Level (RCL), and Cell Level.
Figure 6: An Example of Three Level Table Summarization
Figure 6 shows three summarization strategies applied to the same table. As illustrated, table-level summaries offer an overview but lose fine-grained structural information. The over-abstraction reduces the discriminative power during retrieval, leading to decreased precision and increased ambiguity when distinguishing between similar tables. In addition, cell-level summaries provide highly specific and detailed descriptions, capturing cell-level semantics that are often useful for retrieval. However, this fine granularity comes at the cost of redundancy and low retrieval efficiency, as repeated expressions across cells inflate input length and dilute relevance during retrieval. We do not include cell-level summarization in our experimental design due to its high cost and significant computational overhead.
Therefore, we focus on row and column level (RCL) summarization, which strikes a balance between structure preservation and expression compactness by summarizing content at the header level. To further reduce noise, we additionally mask out non-semantic numeric values in table cells. The RCL middle-ground representation retains the hierarchical structure of tables and preserves header-level semantics essential for retrieval. This design enables more effective alignment between table content and question, serving as a foundation for retrieval in heterogeneous documents.
Appendix C Baselines
This section provides detailed descriptions of the baseline methods used for both retrieval and QA tasks.
For the retrieval task, we include several representative methods:
  • Standard RAG: A basic retrieval-augmented generation framework that retrieves documents using dense similarity search with embeddings from pre-trained models such as DPR, without table-specific optimization.
  • LangChain: A semi-structured retrieval pipeline that processes tabular and textual data by summarizing content with LLMs and conducting similarity-based retrieval.
  • SelfRAG [1]: A self-reflective retrieval framework that dynamically retrieves passages and uses reflection tokens to iteratively assess and refine outputs based on the model’s own feedback.
  • Table Retrieval [4]: A method tailored for table retrieval that discovers join relationships and applies mixed-integer programming-based re-ranking to optimize both table-query and table-table relevance.
For the QA task, we consider several widely used reasoning frameworks that have demonstrated strong performance in answering structured and semi-structured questions:
  • Direct: A simple method where the model directly generates an answer without explicit reasoning or intermediate steps.
  • COT [25]: A reasoning-based prompting strategy that guides the model to generate intermediate reasoning steps, improving its ability to solve complex problems.
  • EEDP [19]: A multi-step reasoning framework that decomposes the answer process into explanation, refinement, and decision phases to enhance accuracy.
  • POT [6]: A technique that prompts the LLM to produce a structured program for reasoning, which is then executed externally to yield precise results, especially for numerical problems.
Appendix D Evaluation Metrics
We adopt two standard metrics to evaluate the performance of the retrieval and QA modules:
  • HiT@K: Measures the proportion of questions for which the correct document is retrieved within the top-K candidates. This metric reflects the effectiveness of the retrieval module. Formally,
    $\text{HiT@K}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left[d_{i}^{\ast}\in\text{TopK}(q_{i})\right],$ (11)
    where $N$ is the total number of questions, $q_{i}$ is the $i$-th question, $d_{i}^{\ast}$ is its ground-truth document, and $\mathbb{I}[\cdot]$ is the indicator function.
  • Exact Match (EM): Calculates the percentage of predicted answers that exactly match the ground-truth answers. This metric evaluates the accuracy of the QA module. Formally,
    $\text{EM}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left[a_{i}=\hat{a}_{i}\right],$ (12)
    where $a_{i}$ and $\hat{a}_{i}$ denote the ground-truth and predicted answers for question $q_{i}$, respectively.
Appendix E Algorithm Pseudo Code of MixRAG
Algorithm 1 illustrates the overall design of MixRAG (excluding heterogeneous document representation) in three sequential stages. (1) Ensemble Retrieval: BM25 and vector-based retrieval run in parallel on the query $Q$, producing $n$ candidates from the table-level corpus and $m$ candidates from the row–column-level corpus, which are then merged into an initial set of $K$ candidate documents. (2) LLM-based Retrieval: For the $K$ candidate documents, we first retain the passages most semantically similar to the question and reconstruct them in their original order into “new documents,” then feed all new documents into the LLM to accurately select the target document $D^{*}$. (3) QA Inference via RECAP: In the RECAP stage, the LLM restates the question, extracts evidence, and produces a computation formula for an external calculator to execute; a rule-based mechanism then selects the final answer $A$ from the LLM’s output and the calculator’s result.
Algorithm 1 MixRAG: Retrieval-Augmented Generation Workflow
Data: $Q$: question, $C_{\mathrm{BM25}}$: BM25 corpus, $C_{\mathrm{Embed}}$: embedding corpus, BM25: BM25 retriever, Embed: dense embedding model, LLM: large language model, Calc: external calculator for executing formulas provided by RECAP, $n$: number of BM25 candidates; $m$: number of embedding candidates; $K$: final candidate set size; $\Theta$: chunk-filter threshold
Result: $A$: predicted answer to the question
1:function MixRAG($Q,\,C_{\mathrm{BM25}},\,C_{\mathrm{H\text{-}RCL}},\,\text{BM25},\,\text{Embed},\,\text{LLM},\,\text{Calc},\,n,\,m,\,K,\,\Theta$)
2:  // Stage 1: Ensemble Retrieval
3:  $S_{1}\leftarrow$ BM25Retrieve($\text{BM25},\,C_{\mathrm{BM25}},\,Q,\,n$)
4:  $S_{2}\leftarrow$ EmbedRetrieve($\text{Embed},\,C_{\mathrm{Embed}},\,Q,\,m$)
5:  $S\leftarrow$ MergeCandidates($S_{1},\,S_{2},\,K$)
6:  // Stage 2: LLM-based Retrieval
7:  $D^{*}\leftarrow$ LLMBasedRetrieve($\text{LLM},\,Q,\,S,\,\Theta$)
8:  // Stage 3: QA Inference (RECAP)
9:  $A\leftarrow$ RECAP($\text{LLM},\,\text{Calc},\,Q,\,D^{*}$)
10:  return AA
11:end function
12:function BM25Retrieve($\text{BM25},\,\text{corpus},\,Q,\,n$)
13:  return top-$n$ documents by BM25 score
14:end function
15:function EmbedRetrieve($\text{Embed},\,\text{corpus},\,Q,\,m$)
16:  return top-$m$ documents by $\cos\bigl{(}\text{Embed}(Q),\,\text{Embed}(\text{chunk})\bigr{)}$
17:end function
18:function MergeCandidates($S_{1},\,S_{2},\,K$)
19:  $U\leftarrow S_{1}\cup S_{2}$ $\triangleright$ remove duplicates
20:  return top-$K$ of $U$
21:end function
22:function LLMBasedRetrieve($\text{LLM},\,Q,\,S,\,\Theta$)
23:  $S^{\prime}\leftarrow\emptyset$
24:  for all $D\in S$ do
25:   $M\leftarrow$ SelectTopChunks($Q,\,D,\,\Theta$) $\triangleright$ retain chunks most relevant to the question
26:   $D^{\prime}\leftarrow$ Recombine(M,DM,\,D)
27:   $S^{\prime}\leftarrow S^{\prime}\cup\{D^{\prime}\}$
28:  end for
[…]
30:end function
31:function RECAP($\text{LLM},\,\text{Calc},\,Q,\,D^{*}$)
32:  $r\leftarrow\text{LLM.Restate}(Q)$ $\triangleright$ (1) Analyze Question Type
33:  $e\leftarrow\text{LLM.Extract}(r,\,D^{*})$ $\triangleright$ (2) Extract Relevant Data
34:  $\ell\leftarrow\text{LLM.Compute}(e)$ $\triangleright$ (3) Perform Logical Reasoning
35:  $a_{\mathrm{LLM}}\leftarrow\text{LLM.Answer}(\ell)$ $\triangleright$ (4) Generate LLM Answer
36:  $f\leftarrow\text{LLM.PresentFormula}(\ell)$
37:  $a_{\mathrm{Calc}}\leftarrow\text{Calc.Evaluate}(f)$ $\triangleright$ (5) Generate Calc Answer
38:  return SelectFinal($a_{\mathrm{LLM}},\,a_{\mathrm{Calc}},\,Q$) $\triangleright$ (6) Provide Final Answer
39:end function
Appendix F Hyperparameter Experiment
In this section, we investigate the effects of two key hyperparameters within the Ensemble Retrieval module of MixRAG. These include the number of documents $n$ selected by the BM25 retriever and the number of top documents $m$ selected by the embedding-based retriever.
The final number of chunks $k$, passed into the LLM-based retrieval stage, is constrained by the total number of candidates retrieved, i.e., $k\leq n+m$. In practice, $k$ determines the number of document chunks provided as input to the language model, and thus directly affects both context length and inference quality. Considering the token limitations and reasoning capacity of LLMs, we empirically find that setting $k\approx 100$ strikes a balance between informativeness and model performance.
Based on this constraint, we conduct a grid search over different combinations of $n$ and $m$ such that $n+m=100$, where $k$ denotes the number of unique document chunks after deduplication. We then evaluate the impact of these configurations on retrieval performance.
Table 5: Retrieval performance under different $(n,m)$ configurations with deduplicated document count $k$.
BM25 ($n$) Embedding ($m$) Average of $k$ HiT Rate
70 30 96.73 0.9016
60 40 94.26 0.9344
50 50 93.18 0.9703
40 60 93.62 0.9805
30 70 94.92 0.9487
*HiT Rate: proportion of ground-truths in the $k$ deduplicated documents.
HiT Rate measures the proportion of cases where the ground-truth document appears within the $k$ deduplicated document chunks returned by Ensemble Retrieval, thereby reflecting the overall retrieval precision. According to Table 5, the configuration with $n=40$ and $m=60$ achieves the highest retrieval performance, reaching a HiT Rate of 0.9805 with an average deduplicated document count of 93.62. This setting offers an effective balance between sparse and dense retrieval, leading to more comprehensive and relevant document coverage. We adopt $n=40$ and $m=60$ as the default hyperparameter setting for Ensemble Retrieval in our subsequent experiments.
Appendix G Comparison of RECAP with Various Baselines
The experimental results of RECAP and baseline methods on the DocRAGLib dataset are shown in Table 6. Five LLMs are used in the experiments, with the EM metric as the evaluation criterion. To ensure fairness in evaluation, we strictly constrain the output formats of the LLMs and tailor the EM-based evaluation mechanism to the different answer types in the dataset.
Among the baseline methods, PoT performs the best, mainly due to its program execution logic, which improves the accuracy of calculation-based questions in the DocRAGLib dataset. However, RECAP outperforms PoT in EM across all five LLMs, demonstrating its superiority. RECAP improves the accuracy of calculation-based questions through reasoning via external calculation functions, reducing output instability compared to program-generated methods like PoT. Furthermore, GPT-4o shows superior performance with the RECAP method (EM = 0.6466) compared to the other four models and is ultimately selected for the Multi-Step Reasoning Module in MixRAG.
Table 6: Comparison of RECAP in MixRAG Document-QA Performance with Various Baselines
Models
Method GPT-4o GPT-4o mini Qwen2.5-32B Mistral-Nemo-2407 LLama3.1-7B
Direct 0.3534 0.2355 0.2180 0.1003 0.1579
CoT 0.4687 0.3684 0.4837 0.2506 0.2531
EEDP 0.4737 0.3759 0.4862 0.2757 0.2080
PoT 0.6190 0.5564 0.5589 0.3409 0.2932
RECAP 0.6466 0.5815 0.6416 0.4411 0.4060
Appendix H Evaluation of the Reranking Performance of Different Models
In the LLM-based Retrieval stage, the input to the LLM is the Top k documents obtained after Ensemble Retrieval. At this stage, the LLM can be viewed as a reranking model. After reconstructing the Top k documents, it reorders them based on the similarity between the document content and the question, ultimately selecting the most relevant documents. In this section, we further evaluate the performance differences between the LLM and conventional reranking models in this fine-grained retrieval reranking task.
In order to benchmark the efficacy of our LLM-based reranking against established retrieval paradigms, we evaluate three distinct families of models: (1) Large Language Model (LLM) rerankers, (2) Cross-encoder rerankers, and (3) Bi-encoder embedding rerankers. For the LLM family, we select GPT-4o and its compact variant GPT-4o mini, as well as GPT-3.5 Turbo, due to their demonstrated ability to perform nuanced contextual matching and logical inference over long documents. These models ingest the entire reconstructed top-k candidate set and exploit autoregressive decoding to implicitly compare semantic relevance across documents.
In contrast, the cross-encoder models, including ms-marco-MiniLM-L12-v2, ms-marco-MiniLM-L6-v2, and ms-marco-MiniLM-L2-v2, perform fine-grained token-level interactions via full-attention between query and document, yielding highly calibrated relevance scores at the expense of greater computational cost. Finally, we include the BGE reranker series (e.g., bge-reranker-v2-m3, bge-base-en-v1.5, bge-large-en-v1.5, and bge-base-en) to represent bi-encoder architectures that first encode query and document separately into fixed embeddings and then score relevance via cosine similarity. Together, these three classes span the methodological spectrum from generative inference to cross-attention scoring and pure embedding matching.
The document and question set used for this performance evaluation is a subset of DocRAGLib, consisting of 100 questions and 280 documents, ensuring that each question has a corresponding document. Using the Ensemble Retrieval Module, we retrieve the corresponding Top k (k=100) documents for each question from the 280 documents. At this point, the HiT@100 for all questions is 0.98, meaning that for 98 of the questions, the Top 100 documents contain the unique document corresponding to each question.
In Table 7, we present the evaluation results of different models in the fine-grained retrieval reranking task. The metrics shown in the table include HiT@1, HiT@3, and HiT@5, which measure the models’ ability to correctly rank and further retrieve the top documents from the Top k documents. From the results, it is evident that GPT-4o performs the best across all evaluation metrics, particularly in scenarios requiring high ranking accuracy. Specifically, GPT-4o achieves a HiT@1 score of 0.73, surpassing GPT-4o mini by 15.9% and the best-performing rerank model, ms-marco-MiniLM-L12-v2, by 25.9%.
Table 7: Evaluation of the Reranking Performance of Different Models
Rerank Method Model HiT@1 HiT@3 HiT@5
LLM GPT-4o 0.73 0.79 0.86
GPT-4o mini 0.63 0.70 0.77
GPT-3.5 Turbo 0.51 0.57 0.69
Cross-Encoder ms-marco-MiniLM-L-12-v2 0.56 0.67 0.79
ms-marco-MiniLM-L-6-v2 0.48 0.58 0.65
ms-marco-MiniLM-L-2-v2 0.52 0.57 0.62
BGE-Reranker bge-reranker-v2-m3 0.58 0.66 0.72
bge-base-en-v1.5 0.53 0.55 0.56
bge-large-en-v1.5 0.49 0.54 0.56
bge-base-en 0.37 0.42 0.47
Appendix I Related Work
I.1 RAG
RAG has significantly enhanced language models, particularly in knowledge-intensive tasks involving document data. This section presents existing work from two main aspects: advancements in RAG techniques [3; 22; 21; 13; 28; 12; 1; 8] and its applications across different modalities [2; 10; 17; 18; 11].
Advancements in RAG techniques [3; 22; 21; 13; 28; 12; 1; 8] aim to improve retrieval accuracy, efficiency, and robustness in knowledge-intensive tasks. RETRO [3] combines document chunks retrieved by local similarity with preceding tokens in auto-regressive models but faces challenges with long documents. DAPR [21] addresses this by integrating hybrid retrieval with BM25 and contextualized passage representations for long document passages. Self-RAG [1] enables adaptive retrieval and self-reflection through special tokens, thereby improving factuality and controllability. However, these methods primarily focus on text-based RAG, with limited progress in handling mixed modalities, which pose unique retrieval and reasoning challenges.
Recent advancements in RAG have extended to multimodal data beyond text. TableRAG [5] applies RAG to single-table QA by selecting schema and cell-level evidence for efficient and scalable reasoning. Chen et al. [4] propose a method combining join relationship discovery and mixed-integer programming-based re-ranking for table retrieval. KRAGEN [17] introduces a knowledge graph-enhanced RAG framework that uses graph-based retrieval to improve factual consistency and reasoning in biomedical problems. MRAG-Bench [11] focuses on vision-centric RAG, highlighting cases where visual knowledge outperforms textual information, especially for vision-language models. However, none of these approaches address RAG for the combined modality of tables and text within documents, a gap that MixRAG specifically addresses.
I.2 Document QA
Document QA refers to the task of extracting or generating answers based on the content of a given document. This section reviews existing approaches to Document QA from two primary perspectives: methods leveraging pre-trained language models [7; 33; 9; 30; 29] and methods utilizing LLM [16; 19; 25; 6; 31].
Pre-trained methods are typically fine-tuned on document-question pairs. While HybridQA [7], TAT-QA [33], and HiTab [9] support hybrid or hierarchical data, they lack robust reasoning capabilities. MultiHiertt [30] introduces complex heterogeneous documents, but existing models like MT2Net and NAPG [29] still struggle with multi-step and numerical reasoning.
Recent studies leverage advanced LLMs for hybrid document QA. HRoT [16] improves performance on the MultiHiertt dataset. EEDP [19], along with prompting strategies like CoT [25] and PoT [6], shows promise across various benchmarks. Zhao et al. [31] assess LLMs’ numerical reasoning over hybrid inputs. Despite these advances, LLMs still fall short of human-level performance, underscoring the need for more effective prompting and reasoning strategies.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Abstract and Section 1
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Section 5
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Section 3
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section 4 and Appendix. In addition, we provide access to the code and dataset, which further enhances reproducibility.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our dataset, code and prompts are available in the anonymized link provided at submission.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Section 4 and Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: While we recognize the importance of reporting statistical variability, our experiments involve repeated LLM invocations, and conducting additional trials to estimate variance is expensive in terms of computation.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Section 4
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our paper complies with the NeurIPS Code of Ethics. No sensitive user data or ethically controversial content is involved in the experiments.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A] .
  49. Justification: We do not foresee societal impacts for this work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper does not involve the release of pretrained models, large-scale generative models, or datasets that carry a high risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Section 4
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our paper introduces a new dataset. An anonymized link is provided at submission.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our paper does not involve any crowdsourcing experiments or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our paper does not involve any experiments with human subjects or crowdsourced participants.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Section 4
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.